# MERLIN: Multimodal Embedding Refinement via LLM-based Iterative Navigation for Text-Video Retrieval-Rerank Pipeline

### v1

@@ 3.5 Iterative Embedding Refinement for Reranking
$\theta=\arccos(\mathbf{e}^{\tilde{a}^{r}}\cdot\mathbf{e}^{r-1})$ (5)
@@ 4.1 Setting
To utilize mutlimodal encoders and LLMs without needing private GPUs, we use Google Multimodal Embedding API [^6] https://cloud.google.com/generative-ai-studio for encoding video and text, and the OpenAI GPT-4o API Achiam et al. (2023) [^7] https://chat.openai.com/ for generating questions and answers. These APIs offers comparable performance and reproducibility on benchmarks without private GPUs.
@@ 4.2 Performance on Text-Video Retrieval
The performance of our system is presented in Table 1, demonstrating its efficacy through multiple rounds of feedback learning, reflecting the system’s ability to iteratively to refine and incorporate feedback. Particularly, MERLIN shows significant improvements with each round of feedback: On the MSR-VTT dataset, MERLIN shows improvements from a R@1 of $44.40$ to $78.00$, on the MSVD from $52.39$ to $77.61$, and on ActivityNet from $56.58$ to $68.44$ by the final round.
@@ 4.3 Average Ranking of QA Rounds
In addition to the retrieval performance presented in Table 1, the effectiveness of the iterative query enrichment is further highlighted by examining the average ranking of the target videos across question-answer rounds. This analysis is helpful for understanding how the process enhances the ranking of the target videos. As illustrated Figure 4, the average ranking of the target video consistently improves each consecutive round across all datasets. For instance, on the MSR-VTT dataset, the average ranking significantly improves from $18.57$ in round 00 to $2.5$ by final round. Similar improvements are observed on other datasets, with the average ranking on MSVD improving from $13.84$ to $2.4$, and on ActivityNet from $6$ to $2.6$. This demonstrates the consistent improvement, thereby confirming the effectiveness of MERLIN in reranking through iterative feedback learning.
4.4 The Effect of $\alpha$ in Iterative Embedding Refinement
@@
Specifically, for MSR-VTT, MERLIN got 44.4/67.60/76.20 for recall@1/5/10 at round 00 respectively but ended up with 61.6/81.20/87.00 respectively at round $5$. For MSVD, MERLIN got 52.39/77.16/84.78 for recall@1/5/10 at round 00 respectively but ended up with 56.87/78.51/84.63 respectively at round $5$.
@@ 5 Conclusion
While our results are promising, we acknowledge that we cannot provide a comprehensive guide for adapting MERLIN to different settings, as we have not extensively explored the impact of changing various components. However, the core principle of integrating user feedback to iteratively refine the query embedding appears to be a robust approach, regardless of pipeline components, the specific domain or data modality. Future work could investigate the generalization of MERLIN to other multimedia retrieval tasks and explore the optimal configurations for different scenarios.
Other limitation of our approach lies in the use of a human-simulating LLM agent for answering questions based on static video frames. While this agent aims to mimic the human feedback process, it lacks the capability to grasp temporal information and attributes that require a high-level understanding of motion and dynamics. Since the LLM agent first generates answers based on static images and then aggregates them, it struggles to capture knowledge about direction, speed, and other temporal aspects present in the videos.
Moreover, as most pre-trained video encoders also have shortcomings in effectively modeling temporal capabilities Liu et al. (2024), our video encoder may be affected by this limitation as well. This creates a kind of chicken-and-egg problem, where video encoders can benefit from temporal-rich information only when they can understand temporal information effectively. Conversely, even if the video question-answering module (or similar counterparts) can handle temporal-rich information, if the video encoder does not possess the same capability, it may not benefit from this information. This temporal modeling challenge is a prevalent issue that the community needs to address collectively.
@@ Acknowledgements
This work was partially supported by Deep Learning Research Group AttentionX.
Appendix A Algorithm
1:encoder $f_{\text{enc}}()$, user query $q\in\mathcal{Q}$, video $v\in\mathcal{V}$, total question-answer round $R$, retrieved top-$k$ videos at round $r$ $\hat{v}^{r}$, $i$-th candidate among top-$k$ videos at round $r$ $\hat{v}^{r}_{i}$
2:Encode $\mathbf{e}_{q}=f_{\text{enc}}(q)$ given user query $q$
3:Encode $\mathbf{e}_{v}=f_{\text{enc}}(v)$ given video $v$
5:Initialize message list $m=[]$
6:for $r=1$ to $R$ do
7:     Append metadata of $\hat{v}^{r-1}_{0}$ to $m$
9:     Append $\hat{q}^{r}$ to $m$
10:     Generate frame-level answers $\big{[}\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]}=\mathcal{M}_{\mathsf{answer}}(\hat{q}^{r}:\text{video in mind})$ (Equation 3)
12:     Encode $\mathbf{e}_{A^{r}}=f_{\text{enc}}(A^{r})$
13:     Refine embedding $e=Refine(e_{q},\cdots,e_{A^{r}})$
15:end for
16:return Reranked retrieved videos $\hat{v}^{r}_{k}$
Appendix B Prompt Template
B.1 Prompt for Question Generation Module
@@
In order to get useful information from user, it is critical to ask good question that could draw user intention. As depicted in Table 2, we set top $1$ ranked video as anchor video and prompted GPT-4o to refer to anchor video’s metadata. In our case, we used the video’s caption as metadata. However, we believe that questions could be more diverse if we could use other data such as Automatic Speech Recognition (ASR) captions, the characteristics of the video, and so on. As MERLIN proceeds with the chat with the user (a user-simulating agent), we stacked previous questions and answers and To encourage GPT-4 to generate diverse questions without repeating previous ones..
B.2 Prompts for Human-Simulating Agent
@@
As human-simulating agent has two steps for answering to the question regarding “video in mind”, we have two different settings for each step. This method lacks in understanding direction, speed and other temporal knowledge as we dealt in the Limitation. However, we experimentally showed that our human-simulating agent is helpful in enriching information.
B.2.1 Prompt for Question-Answering Module
@@
As depicted in Table 3, we sampled frame from video for every 1 second. Then we asynchronously inputed sampled frame and question from MERLIN. We prompted GPT-4o to answer in detail about fact and not just answer in “Yes” or “No”. However this question-answering module is the part that takes up large portion of API cost so video may be sample in wider stride to lower the API cost.
B.2.2 Prompt for Aggregation Module
@@
As depicted in Table 4, we aggregate all the answers generated from the question-answering module. We prompted GPT-4o to aggregate multiple answers made with multiple frame at question-answering module and appended an aggregating example.
Appendix C Case Studies
The main objective of MERLIN is to improve the ranking of failure cases where the target video is not among the top-ranked candidates. At the same time, it is important to keep the success case to stay in the top-ranked candidates while MERLIN proceeds chatting with user. Retrieving the target video among the top-ranked candidates indicates that MERLIN consistently reflects user intention during the conversation. To qualitatively verify that MERLIN performs its tasks according to the aforementioned objectives, we reviewed several case studies. We focused on how MERLIN bring the rank of failure cases.
C.1 Case study for ActivityNet
As shown in the Figure 5, initial ranking of target video was $224$ using paired query from dataset. However as MERLIN augmented query using user’s response, rank boosted to $36\rightarrow 14\rightarrow 4\rightarrow 1$ as round proceeds. During the conversation, MERLIN was able to understand that the user was looking for a video about Christmas themes, featuring two people, and involving gift wrapping. It managed to rank the target video on top with augmented information.
C.2 Case study for MSVD
As shown in Figure 6, initial ranking of target video was $154$ using paired query from dataset. However as MERLIN augmented query using user’s response, rank boosted to $14\rightarrow 1\rightarrow 1\rightarrow 1$ as round proceeds. During the conversation, MERLIN was able to understand that the user was looking for a video about NBA All-Star game, broadcasted on TNT and score board telling 74:75. It managed to rank the target video on top with augmented information at early round and manage to keep top rank during multiple rounds.
C.3 Case study for MSR-VTT
As shown in Figure 7, initial ranking of target video was $361$ using paired query from dataset. However as MERLIN augmented query using user’s response, rank boosted to $197\rightarrow 14\rightarrow 1\rightarrow 1$ as round proceeds. During the conversation, MERLIN was able to understand that the detailed feature and gesture of human featured on “video in mind”. It managed to rank the target video on top with augmented information at early round and manage to keep top rank during multiple rounds.
@@ C.3 Case study for MSR-VTT
You are given caption about certain video(anchor video) and query used to retrieve the anchor video. However this video may not be the exact video the I am looking for.
Your role is to ask question about the video I have in mind to get more information about video. You have 5 rounds and you can only ask one question at a time.
Focus on attributes like number of people, color, shape and etc.
@@
This is caption of retrieved video. Read the video captions and ask some question to gain more information to help find out exact video. Some video may not have caption due to API error saying sorry I can’t provide blah blah. Captions for video: {anchor video’s caption}
@@
Question-answering round prompt
@@
Based on answer, here’s caption of reranked video.
@@
Table 2: The instruction and specification for question generation module in MERLIN using GPT-4o. After initial retrieval at round 00, MERLIN generates question with initial prompt using information of anchor video’s caption. After user answers to the question, MERLIN rerank the and generate question using new anchor and question-answering round prompt.
Details about human-simulating agent (question-answering module)
@@
You are a helpful assistant that answer the question with details. Don’t jsut answer in yes or no. Provide more details(about facts) about the image that might help the questioner.
@@
Table 3: The instruction and specification for video question-answering human-simulating agent using GPT-4o (question-answering module).
Details about human-simulating agent (aggregation module)
@@
The VQA model is designed to answer questions based on images. To apply it to videos, frames are uniformly extracted from the video over time, and the model provides an answer for each frame to a given question. This means that for a single question, there will be multiple answers - one for each extracted frame. Your role is to review all of the individual answers and summarize them to provide a final answer to the original question. When making final answer, don’t user unnecessary words like ‘Based on the individual answers provided by the VQA model,’. Just answer to the question.
For example, if the question is “Did a cookie appear in the video?” and the individual answers from the frames are [“No”, “No”, “Yes”, “No”], then since a cookie appeared in the 3rd frame, you should summarize and answer the question as “Yes”. Length of aggregated answer should be around 30~35 words.
@@
VQA Answer: {Answers from question-answering module}
@@
Table 4: The instruction and specification for video question-answering human-simulating agent using GPT-4o (aggregation module).

### v2

@@ 3.4 Iterative Embedding Refinement for Reranking
$\theta=\arccos(\mathbf{e}^{\tilde{a}^{r}}\cdot\mathbf{e}^{r-1}).$ (5)
@@ 4.1 Setting
To utilize multimodal encoders and LLMs without needing private GPUs, we use Google Multimodal Embedding API [^6] https://cloud.google.com/generative-ai-studio for encoding video and text, and the OpenAI GPT-4o API Achiam et al. (2023) [^7] https://chat.openai.com/ for generating questions and answers. These APIs offers comparable performance and reproducibility on benchmarks without private GPUs.
@@ 4.2 Performance on Text-Video Retrieval
The performance of our system is presented in Table 1, demonstrating its efficacy through multiple rounds of feedback learning, reflecting the system’s ability to iteratively refine and incorporate feedback. Particularly, MERLIN shows significant improvements with each round of feedback: On the MSR-VTT dataset, MERLIN shows improvements of R@1 from $44.40$ to $78.00$, on the MSVD from $52.39$ to $77.61$, and on ActivityNet from $56.58$ to $68.44$ by the final round.
@@ 4.3 Average Ranking of QA Rounds
In addition to the retrieval performance presented in Table 1, the effectiveness of the iterative query enrichment is further highlighted by examining the average ranking of the target videos across question answer rounds. This analysis is helpful for understanding how the process enhances the ranking of the target videos. As illustrated Figure 3, the average ranking of the target video consistently improves each consecutive round across all datasets. For instance, on the MSR-VTT dataset, the average ranking significantly improves from $18.57$ in round 00 to $2.5$ by the final round. Similar improvements are observed on other datasets, with the average ranking on MSVD improving from $13.84$ to $2.4$, and on ActivityNet from $6$ to $2.6$. This demonstrates the consistent improvement, thereby confirming the effectiveness of MERLIN in reranking through iterative feedback learning.
5 Ablation Study
Model Rounds MSR-VTT
R@1 R@5 R@10
Final Query Retrieval (FQR) 5 51.40 71.00 78.80
Refined Reranking (RR) 5 53.60 74.40 81.80
MERLIN 5 78.00 94.20 96.80
Table 2: The performance comparison of video retrieval performance on MSR-VTT using R@K between Final Query Retrieval (FQR), Refined Reranking (RR), and MERLIN. It is worth noting that FQR and RR employ the generated query at final round.
The iterative embedding refinement improves retrieval performance.
The results in Table 2 demonstrate the effectiveness of iterative embedding refinement in improving the retrieval performance. Final Query Retrieval (FQR), which direct retrieves videos using the generate query, achieves a R@1 of $51.40$. Refined Reranking (RR), which applies reranking to the top-100 initial results, improves performance to $53.60$ at R@1. However, MERLIN, which leverages iterative refinement through multiple rounds of interaction between the query and video embeddings, significantly outperforms both methods, reaching a R@1 of $78.00$, demonstrating the advantage of iterative refinement for aligning query representations with video content. The consistent improvements at R@5 and R@10 further highlight the robustness of MERLIN in video retrieval tasks.
The higher $\alpha$ could mitigate the query drift.
@@ 5 Ablation Study
Specifically, for MSR-VTT, MERLIN got $44.4$/$67.60$/$76.20$ for R@1/5/10 at round 00 respectively but ended up with $61.6$/$81.20$/$87.00$ respectively at round $5$. For MSVD, MERLIN got $52.39$/$77.16$/$84.78$ for R@1/5/10 at round 00 respectively but ended up with $56.87$/$78.51$/$84.63$ respectively at round $5$.
6 Case Study
The main objective of MERLIN is to improve the ranking of failure cases where the target video is not among the top-ranked candidates. At the same time, it is important to keep the success case to stay in the top-ranked candidates while MERLIN proceeds to chat with the user. Retrieving the target video among the top-ranked candidates indicates that MERLIN consistently reflects user intention during the conversation. To qualitatively verify that MERLIN performs its tasks according to the aforementioned objectives, we reviewed several case studies. We focused on how MERLIN brings the rank of failure cases.
Case study for ActivityNet
As shown in Figure 4, the initial ranking of the target video was $224$ using a paired query from the dataset. However as MERLIN augmented the query using the user’s response, the rank boosted to $36\rightarrow 14\rightarrow 4\rightarrow 1$ as the round proceeded. During the conversation, MERLIN was able to understand that the user was looking for a video about Christmas themes, featuring two people, and involving gift wrapping. It managed to rank the target video on top with the augmented information.
Case study for MSVD
As shown in Figure 5, the initial ranking of the target video was $154$ using a paired query from the dataset. However as MERLIN augmented the query using the user’s response, the rank boosted to $14\rightarrow 1\rightarrow 1\rightarrow 1$ as the round proceeded. During the conversation, MERLIN was able to understand that the user was looking for a video about the NBA All-Star game, broadcasted on TNT and the scoreboard telling 74:75. It managed to rank the target video on top with the augmented information at an early round and managed to keep the top rank during multiple rounds.
Case study for MSR-VTT
As shown in Figure 6, the initial ranking of the target video was $361$ using a paired query from the dataset. However as MERLIN augmented the query using the user’s response, the rank boosted to $197\rightarrow 14\rightarrow 1\rightarrow 1$ as the round proceeded. During the conversation, MERLIN was able to understand the detailed features and gestures of humans featured on “video in mind”. It managed to rank the target video on top with the augmented information at an early round and managed to keep the top rank during multiple rounds.
@@ 7 Conclusion
While our results are promising, we acknowledge that we cannot provide a comprehensive guide for adapting MERLIN to different settings, as we have not extensively explored the impact of changing various components. However, the core principle of integrating user feedback to iteratively refine the query embedding appears to be a robust approach, regardless of pipeline components, the specific domain, or data modality. Future work could investigate the generalization of MERLIN to other multimedia retrieval tasks and explore the optimal configurations for different scenarios.
Another limitation of our approach lies in the use of a human-simulating LLM agent for answering questions based on static video frames. While this agent aims to mimic the human feedback process, it lacks the capability to grasp temporal information and attributes that require a high-level understanding of motion and dynamics. Since the LLM agent first generates answers based on static images and then aggregates them, it struggles to capture knowledge about direction, speed, and other temporal aspects present in the videos.
Moreover, as most pre-trained video encoders also have shortcomings in effectively modeling temporal capabilities Liu et al. (2024), our video encoder may be affected by this limitation as well. This creates a kind of chicken-and-egg problem, where video encoders can benefit from temporal-rich information only when they can understand temporal information effectively. Conversely, even if the video question answering module (or similar counterparts) can handle temporal-rich information, if the video encoder does not possess the same capability, it may not benefit from this information. This temporal modeling challenge is a prevalent issue that the community needs to address collectively.
@@ Acknowledgements
Nojun Kwak was supported by NRF grant (2021R1A2C3006659) and IITP grants (RS-2022-II220320, RS-2021-II211343), all funded by MSIT of the Korean Government. This work was partially supported by Deep Learning Research Group AttentionX.
Appendix A Prompt Template
A.1 Prompt for Question Generation Module
@@
To get useful information from a user, it is critical to ask good questions that could elicit the user’s intention. As depicted in Table 3, we set the top $1$ ranked video as the anchor video and prompted GPT-4o to refer to the anchor video’s metadata. In our case, we used the video’s caption as metadata. However, we believe that questions could be more diverse if we could use other data such as Automatic Speech Recognition (ASR) captions, the characteristics of the video, and so on. As MERLIN proceeds with the chat with the user (a user-simulating agent), we stacked previous questions and answers and encouraged GPT-4 to generate diverse questions without repeating previous ones.
A.2 Prompts for Human-Simulating Agent
@@
As a human-simulating agent has two steps for answering the question regarding “video in mind”, we have two different settings for each step. This method lacks in understanding direction, speed, and other temporal knowledge as we discussed in Limitation. However, we experimentally showed that our human-simulating agent helps enrich information.
A.2.1 Prompt for Question Answering Module
@@
As depicted in Table 4, we sampled frames from a video for every 1 second. Then we asynchronously input the sampled frames and the question from MERLIN. We prompted GPT-4o to answer in detail about facts and not just answer with “Yes” or “No”. However, this question answering module is the part that takes up a large portion of API cost so the video may be sampled in a wider stride to lower the API cost.
A.2.2 Prompt for Aggregation Module
@@
As depicted in Table 5, we aggregate all the answers generated from the question answering module. We prompted GPT-4o to aggregate multiple answers made with multiple frames at the question answering module and appended an aggregating example.
@@
You are given a caption about a certain video(anchor video) and a query used to retrieve the anchor video. However, this video may not be the exact video that I am looking for.
Your role is to ask questions about the video I have in mind to get more information about the video. You have 5 rounds and you can only ask one question at a time.
Focus on attributes like the number of people, color, shape etc.
@@
This is the caption of the retrieved video. Read the video captions and ask some questions to gain more information to help find out the exact video. Some videos may not have a caption due to an API error saying sorry I can’t provide blah blah. Captions for video: {anchor video’s caption}
@@
Question answering round prompt
@@
Based on the answer, here’s the caption of the reranked video.
@@
Table 3: The instruction and specification for the question generation module in MERLIN using GPT-4o. After initial retrieval at round 00, MERLIN generates a question with an initial prompt using the information of the anchor video’s caption. After the user answers the question, MERLIN reranks the and generates a question using a new anchor and question answering round prompt.
Details about human-simulating agent (question answering module)
@@
You are a helpful assistant that answers the question with details. Don’t just answer in yes or no. Provide more details(about facts) about the image that might help the questioner.
@@
Table 4: The instruction and specification for video question answering human-simulating agent using GPT-4o (question answering module).
Details about the human-simulating agent (aggregation module)
@@
The VQA model is designed to answer questions based on images. To apply it to videos, frames are uniformly extracted from the video over time, and the model provides an answer for each frame to a given question. This means that for a single question, there will be multiple answers - one for each extracted frame. Your role is to review all of the individual answers and summarize them to provide a final answer to the original question. When making the final answer, don’t use unnecessary words like ‘Based on the individual answers provided by the VQA model,’. Just answer the question.
For example, if the question is “Did a cookie appear in the video?” and the individual answers from the frames are [“No”, “No”, “Yes”, “No”], then since a cookie appeared in the 3rd frame, you should summarize and answer the question as “Yes”. The length of the aggregated answer should be around 30~35 words.
@@
VQA Answer: {Answers from question answering module}
@@
Table 5: The instruction and specification for video question answering human-simulating agent using GPT-4o (aggregation module).